One 55-gran serving of a particular cereal supplies 270 mg of sodium, 11% of the recommended daily allowance. How many moles and atoms of sodium are in the recommended daily allowance?
1
Expert's answer
2020-02-27T09:05:07-0500

270 mg*1g/1000mg*1mol/23g=0.0117 mol - 11%

100% - 0.107 mol recommended daily allowance (0.107mol*6.02*10^23molecules/1 mol=0.642*10^23 molecules)